OpenStudy (anonymous):

which of the following is a quantity that could be represented by the number 2*10^-3

OpenStudy (unklerhaukus):

\[2\times10^{-3}=2\times\frac1{10^3}=2\times\frac1{10}\times\frac1{10}\times\frac1{10}\]